# ChaosBench: A Multi-Channel, Physics-Based Benchmark for Subseasonal-to-Seasonal Climate Prediction

### v3

@@ Appendix A Accountability and Reproducibility Statement
Documentation: Finally, our main webpage will keep track of all important updates and latest documentation, and is accessible at https://leap-stc.github.io/ChaosBench/.
@@ B.3 Evaluation
Third, if we have a probabilistic model that generates ensemble forecasts and are supposed to be evaluated with additional probabilistic metrics, we can run:
@@
# --model_name ’vae_s2s’
@@
# --version_num 0
@@
\$ python eval_ensemble.py --model_name <str> --eval_years <int> --version_num <int> --lra5 [...] --oras5 [...]
@@ Appendix C Related Work
On Input Variables. The number of input channels indicates the number of unique variables used for training data-driven models. For instance, in the case of SubseasonalClimateUSA, these include tmin, tmax, tmean, precip_agg, precip_mean, SST, SIC, z-10, z100, z500, z850, u-250, u-925, v-250, v-925, surface_P, RH, SSP, precipitable water, PE, DEM, KG, MJO-phase, MJO-amp, ENSO-I, despite them having similar variable across data sources.
On Target Variables and Agencies. Similarly, the number of target channels represent the variables these benchmarks are aiming for. This is closely related to the number of benchmark agencies, which refers to the number of physics-based simulations used as target, rather than inputs. In the case for SubseasonalClimateUSA, for instance, the number of target channels correspond to two: precipitation and surface temperature, while the number of benchmark agencies is also two: CFSv2 and ECMWF, despite them using multiple other simulations generated from agencies but as inputs; though evaluated on all in their follow-up work [43] despite not initially described in the dataset paper.
@@
On Probabilistic Metrics. The flag for probabilistic metrics indicates whether these benchmarks incorporate probabilistic, in addition to deterministic metrics.
@@ D.1 Observations from Reanalysis Products
D.1.1 LRA5
@@
Table S1: List of LRA5 reanlysis variables
@@
D.1.2 ORAS5
Table S2: List of ORAS5 reanlysis variables
Acronyms Long Name Units
iicethic sea ice thickness m
iicevelu sea ice zonal velocity $ms^{-1}$
iicevelv sea ice meridional velocity $ms^{-1}$
ileadfra sea ice concentration (0-1)
so14chgt depth of 14 isotherm m
so17chgt depth of 17 isotherm m
so20chgt depth of 20 isotherm m
so26chgt depth of 26 isotherm m
so28chgt depth of 28 isotherm m
sohefldo net downward heat flux $Wm^{-2}$
sohtc300 heat content at upper 300m $Jm^{-2}$
sohtc700 heat content at upper 700m $Jm^{-2}$
sohtcbtm heat content for total water column $Jm^{-2}$
sometauy meridonial wind stress $Nm^{-2}$
somxl010 mixed layer depth 0.01 m
somxl030 mixed layer depth 0.03 m
sosaline salinity PSU
sossheig sea surface height m
sosstsst sea surface temperature ${}^{\circ}C$
sowaflup net upward water flux $kg/m^{2}/s$
sozotaux zonal wind stress $Nm^{-2}$
@@ D.2.1 The UK Meteorological Office (UKMO) [30]
National Centers for Environmental Prediction (NCEP) [54]
@@
D.2.2 China Meteorological Administration (CMA) [32]
@@
D.2.3 European Center for Medium-Range Weather Forecasts (ECMWF) [33]
@@ Appendix G Extended Results
(b) Bias ($\rightarrow 0$ is better)
@@
Figure S2: Evaluation results between baseline climatology (black lines) and physics-based control (deterministic) forecasts. At longer forecasting horizon, most physics-based deterministic forecasts perform worse than climatology while maintaining structures as evidenced from their low SpecDiv (barring NCEP). This can suggest knowledge gap, such as poorly represented subgrid phenomena.
(a) RMSE ($\downarrow$ is better)
(b) Bias ($\rightarrow 0$ is better)
(c) MS-SSIM ($\uparrow$ is better)
(d) SpecDiv ($\downarrow$ is better)
Figure S3: Evaluation results between baseline climatology (black lines) and data-driven models including PanguWeather (PW), FourCastNetV2 (FCN2), and GraphCast (GC). Overall, we observe that data-driven models perform significantly worse than climatology on S2S timescale. They also perform poorly on physics-based metrics indicating the lack of predictive power on multi-scale structures. Note: FCN2 lacks q-700 and climatology naturally has low SpecDiv (direct observations).
(a) RMSE ($\downarrow$ is better)
(b) MS-SSIM ($\uparrow$ is better)
(c) SpecDiv ($\downarrow$ is better)
Figure S4: Evaluation results between baseline climatology (black lines) and physics-based ensembles from ECMWF, CMA, UKMO, NCEP. Overall, we observe that ensemble forecasts perform better than their deterministic counterparts.
@@
(c) SpecDiv: ensemble improves deterministic forecasts if $\texttt{ratio}<1$
Figure S5: Metrics ratio e.g., $\texttt{RMSE}_{ens}/\texttt{RMSE}_{det}$ between ensemble and deterministic forecasts, where the former improves the latter by accounting for IC uncertainty that can lead to long-range instability and trajectory divergences.
@@
Figure S6: Probabilistic evaluation on ensemble forecasts indicating current skill limits of 15 days.
@@ G.1 Effects of Different Autoregressive Training Steps; lead_time
Figure S7: Ablation results for incorporating temporal information in an autoregressive scheme for long-range forecast using UNet models. The x-axis represents the number of forecasting days for T850, Z500, Q700 representative tasks. Blue and orange lines illustrate autoregressive scheme with $S=1$ and $S=5$ respectively. Overall we observe that incorporating temporal information improve the vision-based metrics even at longer forecasting timesteps. However, the converse trend is true where incorporating temporal context makes S2S forecast worse off in some physics-based scores.
@@ G.2 Effects of Subset Optimization; headline_vars
G.3 Qualitative Evaluation

### v5

@@ Appendix A Accountability and Reproducibility Statement
Documentation: Finally, our main webpage will keep track of all important updates and latest documentation, and is accessible at https://leap-stc.github.io/ChaosBench.
@@ B.1 Data Preparation
# Optional: SoTa (deterministic) forecasts
\$ ./process.sh panguweather
\$ ./process.sh graphcast
\$ ./process.sh fourcastnetv2
@@ B.3 Evaluation
Third, if we have a probabilistic model that generates ensemble forecasts (e.g., one checkpoint represents one ensemble member) and are supposed to be evaluated with additional probabilistic metrics, we can run:
@@
# --model_name ’unet_ensemble_s2s’
@@
# --version_nums 0 1 2 ## One ensemble member per version
@@
\$ python eval_ensemble.py --model_name <str> --eval_years <int> --version_nums [...] --lra5 [...] --oras5 [...]
B.4 Optional: Processing Multi-Resolution Input
We open-source the data processing script to allow users to process the inputs given different resolution (highest is $0.25$-degree):
# Process inputs with e.g., 0.25-degree resolution
\$ python scripts/process_atmos.py --resolution 0.25 # ERA5
\$ python scripts/process_ocean.py --resolution 0.25 # ORAS5
\$ python scripts/process_land.py --resolution 0.25 # LRA5
@@ Appendix C Related Work
On Input Variables. The number of input channels indicates the number of unique variables used for training data-driven models. For instance, in the case of SubseasonalClimateUSA, these include tmin, tmax, tmean, precip_agg, precip_mean, SST, SIC, z-10, z100, z500, z850, u-250, u-925, v-250, v-925, surface_P, RH, SSP, precipitable water, PE, DEM, KG, MJO-phase, MJO-amp, ENSO-I, despite them having similar (25) variables across data sources.
On Target Variables and Agencies. Similarly, the number of target channels represent the variables these benchmarks are aiming for. This is closely related to the number of benchmark agencies, which refers to the number of physics-based simulations used as target, rather than inputs. In the case for SubseasonalClimateUSA, for instance, the number of target channels correspond to two: precipitation and surface temperature, while the number of benchmark agencies is also two: CFSv2 (NCEP) and IFS (ECMWF), despite them using multiple other simulations generated from agencies but as inputs; though evaluated on all in their follow-up work [50] despite not initially described in the dataset paper.
@@
On Probabilistic Metrics. The flag for probabilistic metrics indicates whether these benchmarks incorporate probabilistic (e.g., CRPS, CRPSS, Spread, SSR), in addition to deterministic metrics.
@@ D.1 Observations from Reanalysis Products
D.1.1 ERA5
The following table indicates the 48 variables that are inferred by physics-based models. Note that the Input ERA5 observations contains ALL fields, including the unchecked boxes:
Table S1: List of ERA5 reanalysis variables
Parameters/Levels (hPa) 1000 925 850 700 500 300 200 100 50 10
Geopotential height, z ($gpm$)
Specific humidity, q ($kg\,kg^{-1}$)
Temperature, t ($K$)
U component of wind, u ($ms^{-1}$)
V component of wind, v ($ms^{-1}$)
Vertical velocity, w ($Pas^{-1}$)
D.1.2 ORAS5
Table S2: List of ORAS5 reanalysis variables
Acronyms Long Name Units
iicethic sea ice thickness m
iicevelu sea ice zonal velocity $ms^{-1}$
iicevelv sea ice meridional velocity $ms^{-1}$
ileadfra sea ice concentration (0-1)
so14chgt depth of 14 isotherm m
so17chgt depth of 17 isotherm m
so20chgt depth of 20 isotherm m
so26chgt depth of 26 isotherm m
so28chgt depth of 28 isotherm m
sohefldo net downward heat flux $Wm^{-2}$
sohtc300 heat content at upper 300m $Jm^{-2}$
sohtc700 heat content at upper 700m $Jm^{-2}$
sohtcbtm heat content for total water column $Jm^{-2}$
sometauy meridonial wind stress $Nm^{-2}$
somxl010 mixed layer depth 0.01 m
somxl030 mixed layer depth 0.03 m
sosaline salinity PSU
sossheig sea surface height m
sosstsst sea surface temperature ${}^{\circ}C$
sowaflup net upward water flux $kg/m^{2}/s$
sozotaux zonal wind stress $Nm^{-2}$
D.1.3 LRA5
@@
Table S3: List of LRA5 reanalysis variables
@@ D.2.1 The UK Meteorological Office (UKMO) [33]
D.2.2 National Centers for Environmental Prediction (NCEP) [61]
@@
D.2.3 China Meteorological Administration (CMA) [35]
@@
D.2.4 European Center for Medium-Range Weather Forecasts (ECMWF) [36]
@@ E.6 PanguWeather, FourCastNetV2, GraphCast
For this work, we process the forecasts at biweekly temporal resolution. In the codebase, we provide the script for further flexibility, for instance:
# Process biweekly, 1.5-degree forecasts for the year 2022
## Panguweather
\$ python scripts/process_sota.py --model_name panguweather --years 2022
## Graphcast
\$ python scripts/process_sota.py --model_name graphcast --years 2022
## FourCastNetV2
\$ python scripts/process_sota.py --model_name fourcastnetv2 --years 2022
@@ Appendix G Extended Results
(b) ACC ($\uparrow$ is better)
(c) Bias ($\rightarrow 0$ is better)
(d) MS-SSIM ($\uparrow$ is better)
(e) SpecDiv ($\downarrow$ is better)
Figure S2: Evaluation results between baseline climatology (black line) and physics-based control (deterministic) forecasts. At longer forecasting horizon, most physics-based deterministic forecasts perform worse than climatology while maintaining structures as evidenced from their low SpecDiv (barring NCEP).
(a) RMSE ($\downarrow$ is better)
(b) ACC ($\uparrow$ is better)
(c) Bias ($\rightarrow 0$ is better)
(d) MS-SSIM ($\uparrow$ is better)
(e) SpecDiv ($\downarrow$ is better)
Figure S3: Evaluation results between baseline climatology (black line) and data-driven models including PanguWeather (PW), FourCastNetV2 (FCN2), and GraphCast (GC). Overall, we observe that data-driven models perform significantly worse than climatology on S2S timescale. They also perform poorly on physics-based metrics indicating the lack of predictive power on multi-scale structures. Note: FCN2 lacks q-700 and climatology naturally has low SpecDiv (direct observations).
(a) RMSE ($\downarrow$ is better)
(b) ACC ($\uparrow$ is better)
@@
Figure S4: Evaluation results between baseline climatology (black line) and physics-based ensembles from ECMWF, CMA, UKMO, NCEP. Overall, we observe that ensemble forecasts perform better than their deterministic counterparts.
@@
(b) ACC: ensemble improves deterministic forecasts if $\texttt{ratio}>1$
(c) MS-SSIM: ensemble improves deterministic forecasts if $\texttt{ratio}>1$
(d) SpecDiv: ensemble improves deterministic forecasts if $\texttt{ratio}<1$
Figure S5: Metrics ratio e.g., $\texttt{RMSE}_{ens}/\texttt{RMSE}_{det}$ between ensemble and deterministic forecasts, where the former improves the latter by accounting for IC uncertainty that can lead to long-range instability and trajectory divergences. Note: $n$ represents the number of ensemble members. The ratio for ACC fluctuates as the scalar value approaches 0.
@@
Figure S6: Probabilistic evaluation on ensemble forecasts indicating current skill limits of 15 days. Note: $n$ represents the number of ensemble members.
@@ G.1 Effects of Different Autoregressive Training Steps; lead_time
Figure S7: Ablation results for incorporating temporal information in an autoregressive scheme for long-range forecast using UNet models. The x-axis represents the number of forecasting days for t-850, z-500, q-700 representative tasks. Blue and orange lines illustrate autoregressive scheme with $s=1$ and $s=5$ respectively. Overall we observe that incorporating temporal information improve the vision-based metrics even at longer forecasting timesteps. However, the converse trend is true where incorporating temporal context makes S2S forecast worse off in some physics-based scores.
@@ G.2 Effects of Subset Optimization; headline_vars
G.3 Effects of Ensemble Forecasts
This section provides additional results for data-driven ensemble approach, and follow similar evaluation process as the physics-based counterpart.
(b) ACC: ensemble improves deterministic forecasts if $\texttt{ratio}>1$
(c) MS-SSIM: ensemble improves deterministic forecasts if $\texttt{ratio}>1$
(d) SpecDiv: ensemble improves deterministic forecasts if $\texttt{ratio}<1$
Figure S8: Metrics ratio e.g., $\texttt{RMSE}_{ens}/\texttt{RMSE}_{det}$ between ensemble and deterministic forecasts, where the former improves the latter by accounting for IC uncertainty that can lead to long-range instability and trajectory divergences. Note: $n$ represents the number of ensemble members.
(a) CRPS ($\downarrow$ is better)
(b) CRPSS ($>0$ is better)
(c) Spread
(d) SSR ($>0$ is better)
Figure S9: Probabilistic evaluation on ensemble forecasts. Note: $n$ represents the number of ensemble members.
G.4 Power Spectra
(a) Task 1
(b) Task 2
Figure S10: Power spectra for ViT/ClimaX demonstrating energy decay/divergence especially for high $k$ as lead time grows.
G.5 Qualitative Evaluation